# Position: The Current AI Conference Model is Unsustainable! Diagnosing the Crisis of Centralized AI Conference

### v1

@@ 1 Introduction
$\bullet\blacktriangle\blacklozenge$ Faculty average contribution The average contribution exceeds exponentially, projected to exceed one paper per month by the 2040s. Meaningful exchanges are diminished by fostering a high-pressure environment. Decoupling submission alleviates deadline pressure and encourages deeper, more thoughtful research.
@@
Table 1: Overview of conference challenges and solutions, illustrating the many-to-many relationship between issues and core goals: $\bullet$ Scientific Advancement, $\blacksquare$ Knowledge Dissemination, $\blacktriangle$ Community Building, $\blacklozenge$ Social Contract: DEI.
@@ 3 Methodology for Assessing the Conference Crisis
  • Publication and Productivity Metrics: Analysis of CSRankings data involved tracking total publication volumes and calculating the Contribution per Person. This latter metric was derived by normalizing publication counts against the number of Effective Faculties (an approach adapted from 18) to measure academic productivity pressure.
@@ 4 Challenges in Current AI Conferences
What further complicates the situation is that publication counts are not merely increasing linearly with faculty growth; they are, in fact, also growing exponentially ($R^{2}=0.978,p<0.001$). We quantify per-capita contributions across fields, and define
$\mathrm{Contribution\ per\ Person}=\frac{\textit{Publication Count}}{\textit{Effective Faculties}}$ (1)
where pub_count represents the number of publications. As shown in Figrue 3, in the past decade, AI faculty publication rates have grown superlinearly—far outpacing the relative stability in other domains—reaching over twice the per-person output of non-AI fields and doubling within a decade to exceed 4.5 papers per person annually. If the trend continues, per-capita output is projected to exceed one paper per month by the 2040s. This misalignment with AI conferences’ core missions, particularly knowledge dissemination (as outlined in Section 2), diminishes meaningful exchanges by fostering a high-pressure environment that heightens mental health strains, such as anxiety impeding open collaborations and excessive competition stifling creativity and risk-taking issues we analyze further below.

### v3

@@ 1 Introduction
$\bullet\blacktriangle\blacklozenge$ Faculty average contribution The average contribution is growing exponentially, projected to exceed one paper every 2 months by 2030. Meaningful exchanges are diminished by fostering a high-pressure environment. Decoupling submission alleviates deadline pressure and encourages deeper, more thoughtful research.
@@
Table 1: Overview of conference challenges and solutions, illustrating the many-to-many relationship between issues and core goals: $\bullet$ Scientific Advancement, $\blacksquare$ Knowledge Dissemination, $\blacktriangle$ Community Building, $\blacklozenge$ Social Contract: DEI2.
22footnotetext: In this paper, DEI principles specifically refer to ensuring a diverse, equitable, and inclusive participation environment in academic conferences. For a detailed discussion, see reference 6.
@@ 3 Methodology for Assessing the Conference Crisis
  • Publication and Productivity Metrics: Analysis of CSRankings data involved tracking total publication volumes and calculating the Per-capita Contribution. This latter metric was derived by normalizing publication counts against the number of Effective Faculties (an approach adapted from 18) to measure academic productivity pressure.
@@ 4 Challenges in Current AI Conferences
What further complicates the situation is that publication counts are not merely increasing linearly with faculty growth; they are, in fact, also growing exponentially ($R^{2}=0.978,p<0.001$ under a log-linear regression model). We quantify per-capita contributions across fields through
$\mathrm{Per\text{-}capita\ Contribution}=\frac{\textit{Publication Count}}{\textit{Effective Faculties}}$ (1)
where pub_count represents the number of publications. As shown in Figure 3, in the past decade, the exponential AI faculty publication rates have far outpaced the relative stability in other domains, reaching over twice the per-person output of non-AI fields and doubling within a decade to exceed 4.5 papers per person annually. If the exponential trend continues, output per faculty is projected to exceed one paper every two months (6 papers annually) by 2030 (per-capita contribution 95% prediction interval [6.34, 8.30]). This misalignment with AI conferences’ core missions, particularly knowledge dissemination (as outlined in Section 2), diminishes meaningful exchanges by fostering a high-pressure environment that heightens mental health strains, such as anxiety impeding open collaborations and excessive competition stifling creativity and risk-taking issues we analyze further below.